HIGH RESOLUTION BROADBAND RATIONAL SPECTROSCOPY: GROUND STATE SPECTRUM OF 1-CHLORO-1-FLUOROETHYLENE

\underline{KELLY M. HOTOPP}, CHANDANA KARUNATILAKA, AMANDA J. SHIRAR, GIANA L. STORCK, and BRIAN C. DIAN, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

A high resolution ground state rotational spectrum of 1-chloro-1-fluoroethylene was observed by using Chirped Pulse Fourier Transform Microwave Spectroscopy (CP-FTMW). This free-jet, broadband (11 GHz) technique has substantially faster acquisition times compared to the Balle Flygare cavity method allowing the rotational spectrum of 1-chloro-1-fluoroethylene (7.5-18.5 GHz) to be recorded in approximately 30 minutes at 65 kHz resolution. In this spectrum, over two hundred thirty lines were fit using Pickett SPFIT for the main chlorine isotopes. In addition, the ^{13}C peaks associated with each chlorine isotope were also observed and assigned. The high signal-to-noise ratio across the entire 11 GHz bandwidth facilitated the assignment of the ^{13}C species in natural abundance. The rotational constants and distortion terms for all assigned isotopes were calculated using various ab initio methods and are compared to the measured values. Structural analysis of the molecule will be discussed.